OpenStudy (anonymous):

The age of a father is 2 less than 7 times the age of his son. In 3 years, the sum of their ages will be 52. If the son’s present age is s years, which equation models this situation? (s+3)+[(7s−2)+3]=52 3s+3+3(7s−2)=52 s+(7s−2)=52 (s−3)+(7s−2)−3=52

OpenStudy (stamp):

The age of a father f is 2 less than 7 times the age of his son s. f = 7s - 2 (f + 3) + (s + 3) = 52

OpenStudy (stamp):

Substitute into the second equation the equation we derived from the first sentence

OpenStudy (stamp):

(7s - 2) + 3 + s + 3 = 52

OpenStudy (stamp):

8s + 4 = 52

OpenStudy (stamp):

8s = 48 s = 6

OpenStudy (stamp):

f = 7s - 2, if s = 6 then f = 7(6) - 2 f = 42 - 2 f = 40

OpenStudy (stamp):

In three years, the son s will be 9 and the father f will be 43. s + f = 52 9 + 43 = 52 It checks out